I NEED THE PSEUDOCODE

New Balance = Principal * (1 + rate/T)T

Principal is the opening balance in the account. Rate is the interest rate, and T is the number of times the interest is compounded during a year (T is 4 if the interest is compounded quarterly). Write a program that prompts the user for the principal, the interest rate, and the number of times the interest is compounded. The output must be formatted in the following format:

Interest Rate: 4.25%
Times Compounded: 12
Principal: $1000.00
Interest: $43.34
Amount in Savings: $1043.34
Answer #1

PSEUDOCODE BALANCE()

1. READ THE PRINCIPAL AMOUNT to PRINCIPAL

2. READ THE INTEREST RATE to RATE

3. READ THE TIMES COMPOUNDED to TIME

4. INTEREST =RATE/100

5. INTEREST = INTEREST/TIME

6. INTEREST = 1+INTEREST

7. `INTEREST = POW(INTEREST,TIME)

8. NEW_BALANCE = PRINCIPAL * INTEREST

9. PRINT “INTEREST RATE : “+ RATE

10. PRINT “TIMES COMPOUNDED:” + TIME

11. PRINT “PRINCIPAL: $”+PRINCIPAL

12. PRINT “INTEREST: $” +(NEW_BALANCE – PRINCIPAL)

13. PRINT “AMOUNT IN SAVINGS: $”+ NEW_BALANCE

PROCESS OF EVALUATION

1. PRINCIPAL = 1000

2. RATE = 4.25%

3. TIME = 12

4. INTEREST = 4.25/100 = 0.0425

5. INTEREST = 0.0425/12 = 0.00354166666

6. INTEREST = 1+ INTEREST = 1.003541666

7. INTEREST = POW(1.003541666,12) = 1.0433377163096154

8. NEW_BALANCE = 1000 * 1.0433377163096154 = 1043.33771630961 = 1043.34